the equation y= -1777x+27,153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is, x=0 corresponds to 2000, x=3 corresponds to 2003, x=4 corresponds to 2004, and so on. predicit the number of gun deaths in 2004 and 2007. in what year will the number of gun deaths be 11,160?
2004_____
2007_____
the predicted number of gun deaths will be 11,160 in the year_____

Answer by stanbon(75887)

the equation g(x)= -1777x+27,153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is, x=0 corresponds to 2000, x=3 corresponds to 2003, x=4 corresponds to 2004, and so on.
Predict the number of gun deaths in 2004 and 2007.
In what year will the number of gun deaths be 11,160?
-----------------
2004_
g(4) = -1777*4+27153 = 20045
----------------
2007_____
g(7) = -1777*7+27153 = 14714
----------------
the predicted number of gun deaths will be 11,160 in the year_2009
11,160 = -1777x + 27153
1777x = 15993
x = 9
==============
Cheers,
Stan H.